ompany A just reported free cash flows of $24 million and expects FCF to grow at a constant rate of 5% forever. The company has $100 million in short-term investments, $200 million debt, $50 million preferred stock, and 10 million shares of common stock outstanding. Their cost of debt is 11% and they are located in a country with no corporate taxes. If the expected risk free rate is 2%, the expected return on the S&P 500 is 8%, and their beta is 1.5, what is the price of one share of Company A's common stock today?
Answer #1

Price = _ECE, Re-q - Re = = R + (Rm - Rg Jß 2 + (8-2) 65 (value) Price = 24 (1.05) ol-0.05 = $420 million Price per share = $